OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and has a slope of 5?

OpenStudy (amistre64):

y=5x-5(4)+1

OpenStudy (anonymous):

Start with point-slope form of a line and isolate y. y-1=5(x-4)

OpenStudy (anonymous):

y = 5x - 3 y = 5x - 1 y = 5x + 3 y = 5x - 19

OpenStudy (anonymous):

D, look at amistre64's reply